On a weekend​ outing, Carlos rents a motorboat for 10 hours to travel down the river and back. The rental operator tells him to go for 4 hours​ downstream, leaving 6 hours to return upstream.
If the river current flows at a speed of 5 mph, how fast must Carlos travel in order to return in 10 hours?

Answer by Theo(13342)

rate * time = distance.

let r = rate
time is in hours
distance is in miles



going downstream, the formula becomes (r+5)*4 = distance downstream
going upstream, the formula becomes (r-5) * 6 = distance upstream

the distance downstream is the same as the distance upstream.

therefore (r+5)*4 = (r-5)*6
simplify to get 4r+20 = 6r-30
subtract 4r from both sides of this equation and add 30 to both sides of this equation to get 50 = 2r
solve for r to get r = 25.

his boat needs to travel at 25 miles per hour.

going downstream, the stream adds 5 miles per hour to that for a total of 30 miles per hour for 4 hours to equal a distance of 120 miles.

coming back upstream, the stream subtracts 5 miles per hour from that for a total of 20 miles per hour for 6 hours to equal a total distance of 120 miles.

carlos needs to travel 25 miles per hour to go downstream and come back upstream in 10 hours.

that's what i get.